# Tensegrity Robot Proprioceptive State Estimation with Geometric Constraints

### v1

@@ V-C Constraints
Figure 5: Illustration of geometric constraints, red arrows represent the vectors chained to each rod center, blue arrows represent their cross-product vector. In the left figure, the geometric constraints are satisfied, with the cross-product vector pointing along the robot’s axial direction. Conversely, the right figure demonstrates a switch to the opposite configuration. The endcap color in the circled area follows yellow-green-pink in the left figure, and yellow-pink-green in the right, which demonstrate the difference of the robot configuration.
@@ VI-A Experiments Setup
We conducted a series of experiments with the cable-actuated tensegrity robot[6] in both simulation environments and real-world settings. The tensegrity robot consists of three bars, each measuring a length of 1.45m, 76mm in diameter, and weighing 4.0kg. An IMU (VectorNav VN-100) is attached to the drive unit at the center of the bar with red endcaps, providing linear acceleration and angular velocity data at 200Hz. The inelastic cables are actuated by brushless DC motors, and the encoders (AS5047P) report the cable lengths at 100Hz. Due to the current hardware limitations, onboard contact sensing has not yet been implemented. Instead, we rely on the Vicon motion capture system (MoCAP) for contact vector $\mathbf{c}_{t}$. The MoCaP system also records the ground-truth robot state $\mathbf{x}_{t}$, which is used to evaluate the performance of our state estimator. Future developments will incorporate contact sensors to enable fully onboard state estimation. In the Mujoco simulation, the physical properties of the robot including dimensions, mass, moment of inertia (MoI), and cable spring constants, as well as sensor biases and noises, are characterized to match the real-world tensegrity robot [6] (see Table. I for the parameters).
@@ VI-B Simulation Shape Reconstruction Results
We quantitatively evaluated shape reconstruction by comparing the distances between the estimated cable-connected endcap positions $\left\|\mathbf{q_{t}^{i}}-\mathbf{q_{t}^{j}}\right\|,(i,j)\in\mathrm{E}$ to the ground-truth cable length $l_{t}^{ij}$ from the simulation, as shown in Fig. 6. Without incorporating geometric constraints, significant shape reconstruction errors were observed due to the opposite chirality of the tensegrity structure, as illustrated in Fig. 5 (right). By incorporating the geometric constraints, the RMSE was reduced to 4.63 cm, which is negligible given the robot’s bar length of 1.45m.
@@ VI-C Simulation Pose Estimation Results
Our pose estimation method builds on the DRIFT[13] Invariant EKF framework. We conducted four simulation experiments: move-forward, move-backward, right-turn, and left-turn, each generating a different trajectory on flat ground. At the beginning of each sequence, the robot remained stationary for a few seconds for sensor initialization, such as accelerometer bias calibration. The robot was controlled via motor-actuated cables to perform predefined maneuvers. In the simulation, contacts were detected for the contact kinematic correction step in the Right Invariant EKF. Ground-truth pose, IMU measurements, cable length measurements, and contact vectors were published using ROS in the Mujoco simulation node. The shape reconstruction node subscribed to these inputs and published the robot shape $\mathbf{Q}$, while the Invariant EKF pose estimator estimated the robot pose in real-time, recording the estimated and ground-truth trajectories for later comparison. In this work, we employ evo [37] to align the initial segment of the estimated trajectory with the ground truth trajectory, as the initial pose is unobservable.
Fig. 7 presents the estimated position and orientation compared to the ground truth for the forward movement dataset, demonstrating robust tracking of both position and orientation. The deviations in the X and Y positions suggests the estimator is well-suited for translational movements, although larger discrepancies were observed along the Z-axis. The orientation estimation captures the oscillatory motions and are generally aligned with the ground truth.
@@
For the right-turn dataset, as shown in Fig. 9, the robot demonstrated a longer 15.70-meter trajectory. Fig. 8 shows comparable performance to the forward dataset, despite a final drift of 0.76 meters over a 15.70-meter trajectory. The drift percentage remained low at 4.84% highlighting the estimator’s consistent accuracy even in more complex maneuvers.
@@
Metric Forward Backward Left-turn Right-turn Forward-realworld
Traj. Length 7.50 6.74 18.61 15.70 5.02
Final Drift 0.3275 0.3386 0.8221 0.7610 0.1174
Drift Percentage 4.37% 5.02% 4.42% 4.85% 2.34%
RPE RMSE Trans. 0.0911 0.0936 0.1159 0.1022 0.1927
RPE RMSE Rot. 2.86 3.72 3.94 3.18 10.84
@@ VI-D Real-world Pose Estimation Results
We conducted similar experiments in a real-world environment as in the forward simulation. Six Vicon motion capture cameras were set up, as shown in Fig. 11, with reflective markers installed on the tips of the robot’s endcaps. The MoCAP system tracked the markers to compute the ground-truth robot state and detect endcap-ground contacts, as discussed in Sec. VI-A. The initialization procedures followed those described in Sec. VI-C. The VN-100 IMU and motor controller board within the tensegrity robot were connected to an Nvidia Jetson Nano Orin, with motion capture data synchronized via ROS. All collected data were stored for subsequent analysis.
As shown in Fig. 10, the state estimator performed reasonably well in estimating both position and orientation in real-world experiments. The x and y positions were accurately estimated, while the z position showed larger deviations (0.25m) over the 5-meter trajectory. In terms of orientation, the pitch angle shows discrepancies, especially in the latter part of the experiment. Despite these issues, the overall performance of the estimator remains promising when transitioning from simulation to real-world implementation. However, improvements in vertical position tracking remain necessary.
@@
The root-mean-square-errors (RMSE) of the Relative Pose Error (RPE), reported in drift per meter, are summarized in Table. II. The average trajectory drift percentage (final drift/trajectory length) across all five datasets was 4.20% . Given the absence of prior work on dead-reckoning or odometry for tensegrity robots, we compared our results to state-of-the-art dead-reckoning and odometry algorithms for other robotic platforms, including a final drift of 3.18% in full-sized vehicles [13], 1.65% in legged robots [38], and 2.38% in the industrial robot Husky [13]. Despite the significantly increased structural and kinematic complexity of tensegrity robots, our state estimator’s performance remains comparable to these systems, demonstrating the robustness of our approach even in this challenging context.
We also performed runtime evaluations using a personal laptop with an i5-11400H CPU. Shape reconstruction required an average time of 4.1 ms, while InEKF propagation and contact correction took 11.33 µs and 17.46 µs, respectively. These results highlight the efficiency of the proposed algorithm, which operates at high frequencies and is suitable for onboard computing in untethered tensegrity robots. This is particularly important for enabling autonomous operations in resource-constrained environments.
@@ VII Conclusion
We proposed a proprioceptive contact-aided Right Invariant EKF-based state estimator for tensegrity robots. Our approach utilizes an optimization method to reconstruct the shape of the 3-bar triangular prism tensegrity robot, providing kinematic information to the filter for correction. The estimator can estimate the robot’s pose in real-time in a real-world setting, using only IMU and motor encoder measurements along with contact information, and achieves an average positional drift percentage of 4.20%. This roprioceptive approach eliminates the need for off-board sensor setups, and delivers pose estimates comparable in accuracy to the state estimation of traditional rigid robots.
However, contact detection currently remains off-board due to hardware constraints. In future iterations, the development of onboard contact sensors will enable fully on-board state estimation for the tensegrity robot. Moreover, the current real-world experiments, while promising, are not yet comprehensive enough to fully demonstrate the state estimation capabilities across a wider range of maneuvers such as point turning or crawling, and different terrains like inclines and declines.
Future works will focus on incorporating contact endcap geometry into the reconstructed kinematics to improve estimation accuracy. Furthermore, we plan to explore onboard contact detection methods, such as using contact sensors or data-driven approaches, to further enhance the robot’s autonomous capabilities.
@@ Acknowledgment
The authors would like to thank Hybrid Dynamic Robotics Lab team member Yilin Ma’s help with the experiment environment setup.

### v2

@@ V-C Optimization constraints
Figure 5: Illustration of geometric constraints, red arrows represent the vectors chained to each rod center, blue arrows represent their cross-product vectors. In the left figure, the geometric constraints are satisfied, with the cross-product vector pointing along the robot’s axial direction to the right. Conversely, the right figure demonstrates a switch to the opposite direction. The endcap colors in the circled area follow yellow-green-pink in the left figure and yellow-pink-green in the right, demonstrating the difference in the robot configuration.
@@ VI-A Experiments Setup
We conducted experiments with the cable-actuated tensegrity robot[6] in both simulation and real-world settings. The tensegrity robot consists of three bars, each measuring a length of 1.45m, 76mm in diameter, and weighing 4.0kg. An IMU (VectorNav VN-100) is attached at the center of the bar with red endcaps, providing linear acceleration and angular velocity data at 200Hz. The inelastic cables are actuated by brushless DC motors with an AS5047P encoder, which operates at 100Hz. The motors are controlled to apply a minimum torque, ensuring cables remain under tension and preventing slackness[6]. The cable lengths are determined from the encoder values with zero-slackness assumption. The distance between the endcaps is then calculated by adding the endcap radius to the measured cable lengths. Due to the current hardware limitations, onboard contact sensing has not yet been implemented. Instead, we rely on the Vicon motion capture system (MoCAP) for contact vector $\mathbf{c}_{t}$. In the MuJoCo simulation, the physical properties of the robot, including dimensions, mass, moment of inertia (MoI), and cable spring constants, as well as sensor biases and noises, are characterized to match the real-world tensegrity robot [6] (see Table. I for the parameters).
@@ VI-B Simulation Shape Reconstruction Results
We quantitatively evaluated shape reconstruction by comparing the distances between the estimated cable-connected endcap positions $\left\|\mathbf{q_{t}^{i}}-\mathbf{q_{t}^{j}}\right\|,(i,j)\in\mathrm{E}$ and the ground-truth cable length $l_{t}^{ij}$ from the simulation, as shown in Fig. 6. Without incorporating geometric constraints, significant shape reconstruction errors were observed due to the opposite chirality of the tensegrity structure, as illustrated in Fig. 5 (right). By incorporating geometric constraints, the RMSE was reduced to 4.63 cm, which is negligible given the robot’s bar length of 1.45 m.
@@ VI-C Simulation Pose Estimation Results
Our pose estimation method builds on the DRIFT[13] Invariant EKF framework. We conducted five simulation experiments: move-forward, move-backward, right-turn, left-turn, and valley, each generating a different trajectory. At the beginning of each sequence, the robot remained stationary for a few seconds for sensor initialization, such as accelerometer bias calibration. The robot was controlled via motor-actuated cables to perform predefined maneuvers. In the simulation, contacts were detected for the contact kinematic correction step in the Right Invariant EKF. IMU measurements, cable length measurements, and contact vectors were published using ROS in the MuJoCo simulation node, while the ground-truth poses are logged for results comparison. The shape reconstruction node subscribed to these inputs and published the robot shape $\mathbf{Q}$, while the Invariant EKF pose estimator estimated the robot pose in real-time, recording both the estimated and ground-truth trajectories for later comparison. We employ evo [37] to align the initial segment of the estimated trajectory with the ground truth trajectory, as the initial pose is unobservable.
Fig. 7 presents the estimated position and orientation compared to the ground truth for the forward movement dataset, demonstrating robust tracking of both position and orientation. The deviations in the X and Y positions suggest the estimator is well-suited for translational movements, although larger discrepancies were observed along the Z-axis. The orientation estimation captures the oscillatory motions and is generally aligned with the ground truth.
@@
Figure 8: The estimated position and orientation are compared with ground truth in the valley dataset, which features a 15-degree downhill followed by a 15-degree uphill scenario.
In the right-turn and valley datasets, shown in Fig. 10 and Fig. 8, the robot successfully navigated longer trajectories of 15.70 m and 22.19 m, respectively. Despite a final drift of 0.76 m over the 15.70-meter trajectory in the right-turn dataset, the drift percentage remained low at 4.84%. Similarly, the valley dataset reported a low drift percentage of 4.15%. As illustrated in Fig. 9 and Fig. 8, the estimator demonstrated comparable qualitative performance to the forward dataset, highlighting its consistent accuracy even during more complex maneuvers and on challenging terrain.
@@
Metric Forward Backward Left-turn Right-turn Forward-realworld valley
Traj. Length 7.50 6.74 18.61 15.70 5.02 22.19
Final Drift 0.3275 0.3386 0.8221 0.7610 0.1174 0.92
Drift percentage 4.37% 5.02% 4.42% 4.85% 2.34% 4.15%
RPE RMSE Trans. 0.0911 0.0936 0.1159 0.1022 0.1927 0.0735
RPE RMSE Rot. 2.86 3.72 3.94 3.18 10.84 4.13
@@ VI-D Real-world Pose Estimation Results
We conducted analogous experiments in a real-world environment to those in the forward simulation. Six Vicon motion capture cameras were arranged as shown in Fig. 12, with reflective markers affixed to the tips of the robot’s endcaps. The MoCAP system tracked these markers to compute the ground-truth robot state and detect endcap-ground contacts, as detailed in Sec. VI-A. The VN-100 IMU and motor controller board inside the tensegrity robot were interfaced with an Nvidia Jetson Orin Nano, with motion capture data synchronized via ROS. All collected data were archived for subsequent analysis.
As illustrated in Fig. 11, the state estimator demonstrated satisfactory performance. The x and y positions were estimated with high accuracy, whereas the z position exhibited larger deviations (0.25 m) over the 5-meter trajectory. Regarding orientation, discrepancies were observed in the pitch angle, particularly toward the latter portion of the experiment. Despite these issues, the overall performance of the estimator remains promising in the transition from simulation to real-world implementation. Potential sources of error contributing to the z-axis discrepancies include the contact impact on the IMU sensor, insufficient motion excitation along the z-axis compared to the x-y plane, and drift in estimated contact points. Collectively, these factors appear to reduce the accuracy of the z-axis position estimation.
@@
The root-mean-square-errors (RMSE) of the Relative Pose Error (RPE), reported in drift per meter, are summarized in Table. II. The average trajectory drift percentage (final drift/trajectory length) across all six datasets is 4.20%. Given the absence of prior work on dead-reckoning or odometry for tensegrity robots, we compared our results to state-of-the-art dead-reckoning and odometry algorithms for other robotic platforms, including a final drift of 3.18% in full-sized vehicles [13], 1.65% in legged robots [38], and 2.38% in the industrial robot Husky [13]. Despite the significantly increased structural and kinematic complexity of tensegrity robots, our state estimator’s performance remains comparable to those of these systems.
Runtime evaluations were performed on a personal laptop with i5-11400H CPU. Shape reconstruction required an average time of 4.1 ms, while InEKF propagation and contact correction took 11.33 µs and 17.46 µs, respectively. These results highlight the efficiency of the proposed algorithm enabling onboard computing on untethered tensegrity robots.
@@ VII Conclusion
We proposed a proprioceptive contact-aided Right Invariant EKF-based state estimator for tensegrity robots. Our approach utilizes an optimization-based method to reconstruct the shape of the 3-bar triangular prism tensegrity robot, providing kinematic information to the filter for correction. The estimator can estimate the robot’s pose in real-time in a real-world setting, relying solely on an IMU, motor encoder measurements, and contact information. It achieves an average positional drift percentage of 4.20%.
This proprioceptive approach eliminates the need for off-board sensor setups while delivering pose estimates comparable in accuracy to traditional rigid robots. However, contact detection currently remains off-board due to hardware constraints. Furthermore, the current real-world experiments are limited in scope and do not fully showcase the state estimation capabilities across a broader range of maneuvers, such as point turning or crawling, and varied terrains.
Future work will focus on incorporating contact endcap geometry into the reconstructed kinematics to improve estimation accuracy. Additionally, we plan to investigate onboard contact detection methods, such as implementing soft endcap contact sensors based on gelslim[39, 40] or leveraging a graph neural network-based approach[41], to enable a fully proprioceptive tensegrity robot system.
@@ Acknowledgment
The authors thank Hybrid Dynamic Robotics Lab team member Yilin Ma’s help with the experiment environment setup.